OpenStudy (anonymous):

Solve the system of equations using the elimination method. x + 4y = 0 x = -5 + y

OpenStudy (anonymous):

Minus the y on the 2nd eqution to get x+4y=0 x-y=-5

OpenStudy (anonymous):

On the 2nd equation multiple a - so it will be x+4y=0 -x+y=5

OpenStudy (anonymous):

Then you will get 5y=5 Divide the 5 on both sides to be y=1 Then plug in the 1 on one of the equtions to get x=-4 & y=1

OpenStudy (anonymous):

Sorry, y=-1